# Predicting the Curie temperature in substitutionally disordered alloys using a first-principles based model

### v1

@@ Abstract
keywords:
Magnetism , Curie temperature , Substitutional Disordered Alloys , Density Functional Theory
[…]
[1]organization=Department of Physics, Chemistry and Biology, addressline=Linköping University, postcode=581 83, city=Linköping, country=Sweden
1 Introduction
@@
Materials with a spontaneous macroscopic magnetization lose it at the Curie temperature ($T_{\text{C}}$) due to a transition from a ferro- or ferrimagnetically ordered state to a disordered paramagnetic state. In the search for technologically applicable magnets, knowledge of this temperature is crucial. For magnetic alloys, the effect of different compositions is of high interest as it enables tailoring $T_{\text{C}}$ of a magnetic material. Conducting reliable predictions of $T_{\text{C}}$ in general and with respect to the composition of an alloy in particular, is therefore of great importance. In alloy systems, changes in composition can modify both the number and strength of magnetic exchange interactions, and the magnetic ordering itself. For example, increasing the concentration of a nonmagnetic element reduces the number of magnetic neighbors, weakening the overall magnetic coupling and lowering $T_{\text{C}}$. Conversely, alloying with elements that possess strong magnetic moments—such as transition metals with partially filled $d$ or $f$ orbitals—can enhance the magnetic interactions within an alloy. These elements contribute to stronger exchange interactions, which stabilize magnetic order at higher temperatures, resulting in an increased Curie temperature [1].
@@
The Curie temperature can be measured by means of experiments and derived using theoretical calculations. The experimental approach quickly becomes too time-consuming and costly to apply when exploring a wide range of new magnetic materials. In theoretical calculations, one may, for example, construct a Heisenberg Hamiltonian and derive the exchange interactions between magnetic moments using the method developed by Liechtenstein et al. [2], subsequently determining $T_{\text{C}}$ using Monte Carlo simulations or a mean-field approximation [3, 4, 5, 6]. This method is, however, often impractical for large-scale investigations of systems with different structures and chemistries, as it requires extensive human manual input. Hence, studies employing this method often focus on only one or a few similar systems [7, 8, 3, 9, 5, 6, 10].
@@
To accelerate theoretical predictions of the Curie temperature, machine learning (ML) methods have been introduced in multiple studies [11, 12, 13, 14, 15, 16, 17, 18, 19]. This approach is highly compatible with a high-throughput search of magnetic materials. However, attempts at developing general ML models for predicting the Curie temperature have faced challenges in capturing a consistently physical and continuous variation of $T_{\text{C}}$ as a function of composition [11, 12].
@@
As mentioned previously, when the composition of an alloy changes, so does the number and type of nearest magnetic neighbors, leading to variations in exchange interactions and the stability of the magnetic ground state. Understanding and incorporating the effects of these factors is crucial for predicting how alloying will influence the magnetic properties of materials. To tackle these systems it therefore appears particularly important to employ physics-based predictive models. In a previous work [20], we created a physics-based model, utilizing output from density functional theory (DFT) calculations, and found it to greatly improve the predictions of the substitutionally disordered alloy Fe1-xCox compared to ML models. The model uses the energy difference between the paramagnetic state and the magnetic ground state, the magnetic entropy, and the number of nearest magnetic neighbors. The paramagnetic state is simulated using a supercell implementation of the disordered local moments (DLM) method [21, 22], and the magnetic ground state is determined by a ground-state search method introduced in the work by Ehn et al. [23]. Details of the DLM simulations and ground-state search are described in Sec. 2.3. In the current work, we explore the applicability of this model further by applying it to estimate $T_{\text{C}}$ of the disordered alloys Fe1-xCox, Fe1-xCrx, Fe1-xVx, NixCu1-xMnSb, NixMnSb, Ti1-xCrxN, and Co1-xAlx. The results are compared to the experimental $T_{\text{C}}$. Additionally, the model is applied to Fe1-xTcx to assess its predictive accuracy in a system where experimental data is scarce due to the radioactive nature of technetium. This allows for theoretical insights that may be useful to guide future experimental validation efforts. We also address and analyze some limitations of our model, discussing their underlying causes.
@@
This current investigation shows that the method from Ref. [20] can be applied across a wide range of systems, unlike other theoretical methods that often require hands-on adjustments for each specific case. Overall, it offers a reliable approach for predicting and tailoring the Curie temperature by adjusting the alloy composition.
2 Method
2.1 Model Description
@@
In Ref. [20], four models were developed to predict the Curie temperature of magnetic materials based on output from DFT calculations. The most elaborate model has the form,
@@
where the parameters $A$, $B$, $C$, and $D$ are fitted to the 32 known ferro- and ferrimagnetic materials presented in Ref. [20] giving $A=0.85$. $B=0.69$, $C=0.14$, and $D=124$; $\Delta E=E_{\text{DLM}}-E_{\text{GS}}$ is the energy difference between the DLM state and the magnetic ground state found after the ground-state search (described more in Section 2.3); $NN$ is the number of nearest magnetic neighbors of the magnetic atoms, defined as atoms with magnetic moments larger than 0.75 $\mu_{B}$ in the magnetic ground state; $S^{\text{mag}}$ is the magnetic entropy of the DLM state given as,
@@
None of the alloys investigated in the present work are included among the 32 systems used to obtain the fitting parameters of Eq. (1) and can thus be seen as a critical test of its generality. The parameters $B$ and $C$ are optimized using the Nelder-Mead method [24]. A linear fit to the experimental $T_{\text{C}}$ subsequently gives the parameters $A$ and $D$ [20].
2.2 Rationale Behind the Model Formulation
The motivation for the form of Eq. (1) is given in more detail in Ref. [20], and is briefly outlined here. There are three physical quantities in Eq. (1): $\Delta E$, $S^{\text{mag}}$, and $NN$. Through a mean-field treatment of the magnetic system an estimate of $T_{\text{C}}$ can be expressed in terms of the energy difference between the paramagnetic state and the ferromagnetic ground state [25, 26, 27]. The connection between the magnetic entropy, $S^{\text{mag}}$, and $T_{\text{C}}$ can be argued via the approximation of the free energy per magnetic atom,
@@ 2.2 Rationale Behind the Model Formulation
assuming that the ordered magnetic state remains ideally ordered and the paramagnetic state is ideally disordered, thus neglecting all effects of short-range order (SRO). These SRO effects can be shown to vary in importance depending on whether the magnetic energy is distributed over many or few magnetic nearest neighbors. Based on the findings presented in Ref. [20], the expression for $T_{\text{C}}$ in Eq. (4) can be adjusted with a factor $1-\frac{B}{NN^{C}}$ to capture the different impact of SRO effects in different crystal structures and alloy compositions. This adjustment results in a lower estimated $T_{\text{C}}$ for a given energy difference when the number of nearest neighbors is small. The parameters B and C are the same as in Eq. (1).
@@
2.3 Density Functional Theory Calculations
To obtain the energies of the paramagnetic and ground states, the disordered local moments method and a ground-state search are used, respectively. The ground-state search is performed using the method presented by Ehn et al. [23] and described in Ref. [20]. This approach involves running noncollinear magnetic calculations with initial random directions for the magnetic moments. During ionic relaxation, the directions are free to rotate towards their energetically preferred orientations and adjust in magnitude. To simulate substitutionally disordered alloys, the special quasirandom structure method (SQS) is used [28]. We perform a single ground-state search with these SQS cells and compare the total energy to that of a collinear ferromagnetic configuration. In the investigated cases, the energies were found to be comparable, suggesting that a ferromagnetic ground state is the correct ground state, consistent with experimental expectations for these systems.
The DLM method used here is based on the supercell approach implemented by Alling et al. [21]. In the present work, the magnetic moments are set with random directions (in a noncollinear configuration) and with magnetic moment sizes based on the average sizes of each element from the ground-state search. Only one DLM configuration is used to limit the number of computationally intensive calculations. To ensure a nearly ideal DLM configuration, the directions are set so that the SRO is as close to zero as possible. The SRO is defined as,
@@ 2.3 Density Functional Theory Calculations
where, $N_{\text{at}}$ is the total number of atoms in the cell and $N_{\text{NN}}$ is the number of magnetic neighbors in the first coordination shell. In the DLM calculation, the directions are constrained using the method by Ma and Dudarev [29], where the direction and sign of the magnetic moment are fixed, preventing rotation or flips, while allowing its magnitude to vary. Only elements with magnetic moment magnitudes above 0.75 $\mu_{B}$ in the ground-state search are constrained in the DLM run. This limit must, however, be altered in some of the investigated cases; this will be discussed in Sec. 5. The constraining parameter $\lambda$, which determines the strength of the penalty energy, is set incrementally until the slope, given by
@@ 3 Computational Details
To obtain the energies and the magnetic moment sizes, we use DFT calculations as implemented in the Vienna Ab initio Simulation Package (VASP) [30, 31, 32, 33] using projector augmented-wave potentials (PAW) [34, 35] and the Perdew-Burke-Ernzerhof (PBE) generalized gradient approximation (GGA) for approximating the exchange-correlation functional [36]. The energy cutoff is set to 450 eV. The chemical disorder in the investigated alloys is modeled using the special quasi-random structure (SQS) method [28]. For iron-containing alloys, 64-atom SQS supercells are used, except for compositions with 5% or 10% of non-iron elements, where 125-atom supercells are employed. Supercells for $\mathrm{Ti_{1-x}Cr_{x}N}$ contain 128 atoms, while $\mathrm{Ni_{x}Cu_{1-x}MnSb}$ is modeled using 96-atom supercells. In the case of $\mathrm{Ni_{x}MnSb}$, the supercell size increases with the nickel content. Starting from a 96-atom NiMnSb supercell, additional nickel atoms are added as needed. For example, at $x=1.5$, 16 extra nickel atoms are included, resulting in a total of 112 atoms in the supercell. A Monkhorst-Pack 2$\times$2$\times$2 k-point grid is used for all substitutionally disordered alloys [37].
@@
To simulate the paramagnetic state, the DLM approach is used [21, 22]. The calculations are spin-polarized, utilizing noncollinear magnetism. The magnetic moments are initially set to 3 Bohr magneton ($\mu_{B}$) in random directions in the ground-state search. The magnetic moments are then allowed to relax in both magnitude and direction. In the DLM simulations, the directions of the magnetic moments are constrained using the method presented in Ref. [29]. The magnetic moment magnitudes are allowed to relax in the DLM run, but are initially set to 110% of the relaxed average ground-state size of the magnetic moments of the specific atomic species.
4 Results
4.1 Prediction of $T_{\text{C}}$ Across Compositions
@@
Seven different substitutional random alloy systems are investigated using the model of Eq. (1) and compared with experimental results. Fig. 1 shows the results of predicted $T_{\text{C}}$ as a function of composition for body-centered cubic (bcc) and face-centered cubic (fcc) Fe1-xCox, bcc Fe1-xVx, bcc Fe1-xCrx, NixCu1-xMnSb, NixMnSb, and Ti1-xCrxN. The numerical values are found in Table 1. The different Fe-containing alloys in the upper part of Fig. 1 are mainly of the bcc structure. However, for Fe1-xCox there is a phase transition from bcc to fcc coinciding with the magnetic transition between 31 and 78 at % iron in the cobalt- and iron-rich regions, respectively [38]. Hence, for the cobalt-rich region the random alloy is of fcc structure. The experimental results for the Fe1-xCox system are from Ref. [38], which compiled the data from Refs. [39, 40]. The experimental results for Fe1-xVx are from Ref. [41], which is a collection of results from several experimental studies [42, 43, 44, 45, 46, 47, 48, 49]. The experimental Curie temperatures for the Fe1-xCrx are from the Ref. [50].
@@
Figure 1: Predicted and experimental Curie temperatures as a function of composition are shown for the following alloy systems: bcc and fcc Fe1-xCox (experimental data from [38, 39, 40]), bcc Fe1-xVx (experimental data from [41, 42, 43, 44, 45, 46, 47, 48, 49]), bcc Fe1-xCrx (experimental data from [50]), NixCu1-xMnSb (experimental data from [51]), NixMnSb (experimental data from [52]), and Ti1-xCrxN (experimental data from [53]). Here, $x$ represents the fractional substitution of one element for another in each alloy, except for NixMnSb, where $x$ denotes the Ni concentration, varying from 1 to 2. The numerical values of this figure are found in Table 1.
@@
In the NixMnSb system, $x$ represents the concentration of Ni, ranging from 1 to 2, indicating a variation in the Ni content from half to full Heusler alloys, i.e., NiMnSb to Ni2MnSb. In contrast, in the NixCu1-xMnSb system, $x$ is the fractional substitution of Ni for Cu, ranging from 0 to 1. The structure of the Heusler alloy Ni2MnSb has been found in earlier investigation to be of the L21 type, whereas the half Heusler alloy NiMnSb has a C1b structure [54, 55]. Webster and Mankikar found that from NiMnSb to Ni1.6MnSb the structure is ordered in the C1b structure. However, at higher nickel concentrations, there is significant disorder from this structure [56]. The investigated compositions in Fig. 1 for NixMnSb are all in the C1b structure. The experimental $T_{\text{C}}$ for NixCu1-xMnSb are from the work by Ren et al. [51]. For NixMnSb, the experimental results are taken from the recent work of Neibecker et al. [52].
@@
Ferromagnetism was discovered experimentally in Ti1-xCrxN by Refs. [57, 53] at the TiN-rich regime. In Ref. [6], this system was investigated using first-principles calculations, confirming the experimental findings. The experimental results in Fig. 1 for this system are from Ref. [53]. When conducting the ground-state search, the chromium moments do not end up completely ferromagnetically aligned in any of the cases $x=0.125,0.25,0.375$, or $0.5$. Judging from the total magnetic moment of the supercell, the average magnetic moment is 1.84, 0.79, 0.55, and 1.12 $\mu_{B}$/chromium atom for the respective $x$-values, reflecting directional variations but no change in magnitude. In the ferromagnetic case, the average chromium moment has a magnitude of approximately $2.2$ $\mu_{B}$ for all $x$-values. The energies of the ground-state search are compared to a ferromagnetic calculation, and in all cases the ground-state search results in a slightly lower energy. The largest difference is for $x=0.5$, where the ferromagnetic calculation results in an energy 13 meV/chromium atom higher than the ground-state search.
@@
In the workflow developed in our previous work, the magnetic moments to be constrained in the DLM run are determined based on the average magnetic moment magnitude of the specific species obtained in the ground-state search. If the average magnetic moment magnitude is larger than 0.75 $\mu_{B}$ in the ferromagnetic ground state, the magnetic moments are constrained; otherwise they are free to change size and direction in the DLM calculations. These atoms are essentially treated on an equal footing with nonmagnetic atoms. Following this strictly in some cases is, however, not possible. For Fe1-xCrx, the chromium moments are on average approximately $1.3$ $\mu_{B}$ when $x=0.125$ and should therefore be constrained in the DLM run. These moments are, despite this, not constrained to avoid convergence problems, which is discussed in more detail in Sec. 5. This is also the case for the vanadium moments in Fe1-xVx, where, for all tested $x$-values, these moments are larger than $0.75$ $\mu_{B}$ but are never constrained in the DLM calculations.
@@
The fcc Co1-xAlx system is also investigated using our model. This system is separated from the others as an example of a case with more itinerant type of magnetism to illustrate when our model is not expected to perform as well as in the above examples. This is confirmed in Fig. 2. The procedure of increasing the constraining parameter $\lambda$ is slightly adjusted for this alloy system due to a higher sensitivity in the electronic convergence of the self-consistency cycle. After $\lambda=5$ the $\lambda$ parameter is set to 7. In the case of 25% aluminum, a satisfying level of electronic convergence is not reached for $\lambda=10$. Therefore, Fig. 2 shows the results of $\lambda=7$ for all percentages of aluminum. The penalty energy in these cases is approximately $0.2$ meV/magnetic atom, which we consider an acceptable level of accuracy. The experimental values for Co1-xAlx are from Ref. [58]. In Fig. 3a, the magnetic moment sizes of the cobalt atoms are shown for both Co0.9Al0.1 and Co0.9Fe0.1. The moment sizes are given in the DLM state and in the magnetic ground state for comparison. In Fig. 3b, the magnetic moment sizes of iron and aluminum are given for the same systems.
@@
Figure 2: Predicted and experimental Curie temperatures of fcc Co1-xAlx as a function of the fractional substitution of cobalt with aluminum. The experimental Curie temperatures are from Ref. [58].

### v2

@@ Abstract
preprint: AIP/123-QED
I Introduction
@@
Materials with a spontaneous macroscopic magnetization lose it at the Curie temperature ($T_{\text{C}}$) due to a transition from a ferro- or ferrimagnetically ordered state to a disordered paramagnetic state. In the search for technologically applicable magnets, knowledge of this temperature is crucial. For magnetic alloys, the effect of different compositions is of high interest as it enables tailoring $T_{\text{C}}$ of a magnetic material. Conducting reliable predictions of $T_{\text{C}}$ in general and with respect to the composition of an alloy in particular, is therefore of great importance. In alloy systems, changes in composition can modify both the number and strength of magnetic exchange interactions, and the magnetic ordering itself. For example, increasing the concentration of a nonmagnetic element reduces the number of magnetic neighbors, weakening the overall magnetic coupling and lowering $T_{\text{C}}$. Conversely, alloying with elements that possess strong magnetic moments—such as transition metals with partially filled $d$ or $f$ orbitals—can enhance the magnetic interactions within an alloy. These elements contribute to stronger exchange interactions, which stabilize magnetic order at higher temperatures, resulting in an increased Curie temperature haglund_curie_1982 .
@@
The Curie temperature can be measured by means of experiments and derived using theoretical calculations. The experimental approach quickly becomes too time-consuming and costly to apply when exploring a wide range of new magnetic materials. In theoretical calculations, one may, for example, construct a Heisenberg Hamiltonian and derive the exchange interactions between magnetic moments using the method developed by Liechtenstein et al. liechtenstein_local_1987 , subsequently determining $T_{\text{C}}$ using Monte Carlo simulations or a mean-field approximation rusz_exchange_2006 ; lezaic_first-principles_2007 ; alling_effect_2009 ; alling_theory_2010 . This method is, however, often impractical for large-scale investigations of systems with different structures and chemistries, as it requires extensive human manual input. Hence, studies employing this method often focus on only one or a few similar systems halilov_adiabatic_1998 ; kubler_ab_2006 ; rusz_exchange_2006 ; rusz_ab_2006 ; alling_effect_2009 ; alling_theory_2010 ; zelai_first_2023 .
@@
To accelerate theoretical predictions of the Curie temperature, machine learning (ML) methods have been introduced in multiple studies nelson_predicting_2019 ; belot_machine_2023 ; dam_important_2018 ; long_accelerating_2021 ; lu_–fly_2022 ; nguyen_ensemble_2019 ; xu_predicting_2024 ; hu_searching_2020 ; zhai_accelerated_2018 . This approach is highly compatible with a high-throughput search of magnetic materials. However, attempts at developing general ML models for predicting the Curie temperature have faced challenges in capturing a consistently physical and continuous variation of $T_{\text{C}}$ as a function of composition nelson_predicting_2019 ; belot_machine_2023 .
@@
As mentioned previously, when the composition of an alloy changes, so does the number and type of nearest magnetic neighbors, leading to variations in exchange interactions and the stability of the magnetic ground state. Understanding and incorporating the effects of these factors is crucial for predicting how alloying will influence the magnetic properties of materials. To tackle these systems it therefore appears particularly important to employ physics-based predictive models. In a previous work brannvall_predicting_2024 , we created a physics-based model, utilizing output from density functional theory (DFT) calculations, and found it to greatly improve the predictions of the substitutionally disordered alloy Fe1-xCox compared to ML models. The model uses the energy difference between the paramagnetic state and the magnetic ground state, the magnetic entropy, and the number of nearest magnetic neighbors. The paramagnetic state is simulated using a supercell implementation of the disordered local moments (DLM) method alling_effect_2010 ; gyorffy_first-principles_1985 , and the magnetic ground state is determined by a ground-state search method introduced in the work by Ehn et al. ehn_first-principles_2023 . Details of the DLM simulations and ground-state search are described in Sec. II.3. In the current work, we explore the applicability of this model further by applying it to estimate $T_{\text{C}}$ of the disordered alloys Fe1-xCox, Fe1-xCrx, Fe1-xVx, NixCu1-xMnSb, NixMnSb, Ti1-xCrxN, and Co1-xAlx. The results are compared to the experimental $T_{\text{C}}$. Additionally, the model is applied to Fe1-xTcx to assess its predictive accuracy in a system where experimental data is scarce due to the radioactive nature of technetium. This allows for theoretical insights that may be useful to guide future experimental validation efforts. We also address and analyze some limitations of our model, discussing their underlying causes.
@@
This current investigation shows that the method from Ref. brannvall_predicting_2024 can be applied across a wide range of systems, unlike other theoretical methods that often require hands-on adjustments for each specific case. Overall, it offers a reliable approach for predicting and tailoring the Curie temperature by adjusting the alloy composition.
II Method
II.1 Model Description
@@
In Ref. brannvall_predicting_2024 , four models were developed to predict the Curie temperature of magnetic materials based on output from DFT calculations. The most elaborate model has the form,
@@
where the parameters $A$, $B$, $C$, and $D$ are fitted to the 32 known ferro- and ferrimagnetic materials presented in Ref. brannvall_predicting_2024 giving $A=0.85$. $B=0.69$, $C=0.14$, and $D=124$; $\Delta E=E_{\text{DLM}}-E_{\text{GS}}$ is the energy difference between the DLM state and the magnetic ground state found after the ground-state search (described more in Section II.3); $NN$ is the number of nearest magnetic neighbors of the magnetic atoms, defined as atoms with magnetic moments larger than 0.75 $\mu_{B}$ in the magnetic ground state; $S^{\text{mag}}$ is the magnetic entropy of the DLM state given as,
@@
None of the alloys investigated in the present work are included among the 32 systems used to obtain the fitting parameters of Eq. (1) and can thus be seen as a critical test of its generality. The parameters $B$ and $C$ are optimized using the Nelder-Mead method nelder_simplex_1965 . A linear fit to the experimental $T_{\text{C}}$ subsequently gives the parameters $A$ and $D$ brannvall_predicting_2024 .
II.2 Rationale Behind the Model Formulation
The motivation for the form of Eq. (1) is given in more detail in Ref. brannvall_predicting_2024 , and is briefly outlined here. There are three physical quantities in Eq. (1): $\Delta E$, $S^{\text{mag}}$, and $NN$. Through a mean-field treatment of the magnetic system an estimate of $T_{\text{C}}$ can be expressed in terms of the energy difference between the paramagnetic state and the ferromagnetic ground state turek_ab_2003 ; liechtenstein_magnetic_1983 ; bergqvist_theoretical_2007 . The connection between the magnetic entropy, $S^{\text{mag}}$, and $T_{\text{C}}$ can be argued via the approximation of the free energy per magnetic atom,
@@ II.2 Rationale Behind the Model Formulation
assuming that the ordered magnetic state remains ideally ordered and the paramagnetic state is ideally disordered, thus neglecting all effects of short-range order (SRO). These SRO effects can be shown to vary in importance depending on whether the magnetic energy is distributed over many or few magnetic nearest neighbors. Based on the findings presented in Ref. brannvall_predicting_2024 , the expression for $T_{\text{C}}$ in Eq. (4) can be adjusted with a factor $1-\frac{B}{NN^{C}}$ to capture the different impact of SRO effects in different crystal structures and alloy compositions. This adjustment results in a lower estimated $T_{\text{C}}$ for a given energy difference when the number of nearest neighbors is small. The parameters B and C are the same as in Eq. (1).
@@
II.3 Density Functional Theory Calculations
To obtain the energies of the paramagnetic and ground states, the disordered local moments method and a ground-state search are used, respectively. The ground-state search is performed using the method presented by Ehn et al. ehn_first-principles_2023 and described in Ref. brannvall_predicting_2024 . This approach involves running noncollinear magnetic calculations with initial random directions for the magnetic moments. During ionic relaxation, the directions are free to rotate towards their energetically preferred orientations and adjust in magnitude. To simulate substitutionally disordered alloys, the special quasirandom structure method (SQS) is used zunger_special_1990 . We perform a single ground-state search with these SQS cells and compare the total energy to that of a collinear ferromagnetic configuration. In the investigated cases, the energies were found to be comparable, suggesting that a ferromagnetic ground state is the correct ground state, consistent with experimental expectations for these systems.
The DLM method used here is based on the supercell approach implemented by Alling et al. alling_effect_2010 . In the present work, the magnetic moments are set with random directions (in a noncollinear configuration) and with magnetic moment sizes based on the average sizes of each element from the ground-state search. Only one DLM configuration is used to limit the number of computationally intensive calculations. To ensure a nearly ideal DLM configuration, the directions are set so that the SRO is as close to zero as possible. The SRO is defined as,
@@ II.3 Density Functional Theory Calculations
where, $N_{\text{at}}$ is the total number of atoms in the cell and $N_{\text{NN}}$ is the number of magnetic neighbors in the first coordination shell. In the DLM calculation, the directions are constrained using the method by Ma and Dudarev ma_constrained_2015 , where the direction and sign of the magnetic moment are fixed, preventing rotation or flips, while allowing its magnitude to vary. Only elements with magnetic moment magnitudes above 0.75 $\mu_{B}$ in the ground-state search are constrained in the DLM run. This limit must, however, be altered in some of the investigated cases; this will be discussed in Sec. V. The constraining parameter $\lambda$, which determines the strength of the penalty energy, is set incrementally until the slope, given by
@@ III Computational Details
To obtain the energies and the magnetic moment sizes, we use DFT calculations as implemented in the Vienna Ab initio Simulation Package (VASP) kresse_ab_1993 ; kresse_ab_1994 ; kresse_efficiency_1996 ; kresse_efficient_1996 using projector augmented-wave potentials (PAW) blochl_projector_1994 ; kresse_ultrasoft_1999 and the Perdew-Burke-Ernzerhof (PBE) generalized gradient approximation (GGA) for approximating the exchange-correlation functional perdew_generalized_1996 . The energy cutoff is set to 450 eV. The chemical disorder in the investigated alloys is modeled using the special quasi-random structure (SQS) method zunger_special_1990 . For iron-containing alloys, 64-atom SQS supercells are used, except for compositions with 5% or 10% of non-iron elements, where 125-atom supercells are employed. Supercells for $\mathrm{Ti_{1-x}Cr_{x}N}$ contain 128 atoms, while $\mathrm{Ni_{x}Cu_{1-x}MnSb}$ is modeled using 96-atom supercells. In the case of $\mathrm{Ni_{x}MnSb}$, the supercell size increases with the nickel content. Starting from a 96-atom NiMnSb supercell, additional nickel atoms are added as needed. For example, at $x=1.5$, 16 extra nickel atoms are included, resulting in a total of 112 atoms in the supercell. A Monkhorst-Pack 2$\times$2$\times$2 k-point grid is used for all substitutionally disordered alloys monkhorst_special_1976 .
@@
To simulate the paramagnetic state, the DLM approach is used alling_effect_2010 ; gyorffy_first-principles_1985 . The calculations are spin-polarized, utilizing noncollinear magnetism. The magnetic moments are initially set to 3 Bohr magneton ($\mu_{B}$) in random directions in the ground-state search. The magnetic moments are then allowed to relax in both magnitude and direction. In the DLM simulations, the directions of the magnetic moments are constrained using the method presented in Ref. ma_constrained_2015 . The magnetic moment magnitudes are allowed to relax in the DLM run, but are initially set to 110% of the relaxed average ground-state size of the magnetic moments of the specific atomic species.
IV Results
[…]
@@ IV Results
Seven different substitutional random alloy systems are investigated using the model of Eq. (1) and compared with experimental results. Fig. 1 shows the results of predicted $T_{\text{C}}$ as a function of composition for body-centered cubic (bcc) and face-centered cubic (fcc) Fe1-xCox, bcc Fe1-xVx, bcc Fe1-xCrx, NixCu1-xMnSb, NixMnSb, and Ti1-xCrxN. The numerical values are found in Table 1. The different Fe-containing alloys in the upper part of Fig. 1 are mainly of the bcc structure. However, for Fe1-xCox there is a phase transition from bcc to fcc coinciding with the magnetic transition between 31 and 78 at % iron in the cobalt- and iron-rich regions, respectively nishizawa_cofe_1984 . Hence, for the cobalt-rich region the random alloy is of fcc structure. The experimental results for the Fe1-xCox system are from Ref. nishizawa_cofe_1984 , which compiled the data from Refs. 69Stu ; 75Nor . The experimental results for Fe1-xVx are from Ref. smith_fev_1984 , which is a collection of results from several experimental studies 09Por ; 10Hon ; 25Mau ; 30Oya ; 30Wev ; 31Vog ; 34Abr ; 36Fal . The experimental Curie temperatures for the Fe1-xCrx are from the Ref. wijn_3d_1997 .
@@
Figure 1: Predicted and experimental Curie temperatures as a function of composition are shown for the following alloy systems: bcc and fcc Fe1-xCox (experimental data from nishizawa_cofe_1984 ; 69Stu ; 75Nor ), bcc Fe1-xVx (experimental data from smith_fev_1984 ; 09Por ; 10Hon ; 25Mau ; 30Oya ; 30Wev ; 31Vog ; 34Abr ; 36Fal ), bcc Fe1-xCrx (experimental data from wijn_3d_1997 ), NixCu1-xMnSb (experimental data from ren_magnetic_2005 ), NixMnSb (experimental data from neibecker_atomic_2022 ), and Ti1-xCrxN (experimental data from inumaru_ferromagnetic_2007 ). Here, $x$ represents the fractional substitution of one element for another in each alloy, except for NixMnSb, where $x$ denotes the Ni concentration, varying from 1 to 2. The numerical values of this figure are found in Table 1.
@@
In the NixMnSb system, $x$ represents the concentration of Ni, ranging from 1 to 2, indicating a variation in the Ni content from half to full Heusler alloys, i.e., NiMnSb to Ni2MnSb. In contrast, in the NixCu1-xMnSb system, $x$ is the fractional substitution of Ni for Cu, ranging from 0 to 1. The structure of the Heusler alloy Ni2MnSb has been found in earlier investigation to be of the L21 type, whereas the half Heusler alloy NiMnSb has a C1b structure webster_magnetic_1968 ; szytula_atomic_1972 . Webster and Mankikar found that from NiMnSb to Ni1.6MnSb the structure is ordered in the C1b structure. However, at higher nickel concentrations, there is significant disorder from this structure webster_chemical_1984 . The investigated compositions in Fig. 1 for NixMnSb are all in the C1b structure. The experimental $T_{\text{C}}$ for NixCu1-xMnSb are from the work by Ren et al. ren_magnetic_2005 . For NixMnSb, the experimental results are taken from the recent work of Neibecker et al. neibecker_atomic_2022 .
@@
Ferromagnetism was discovered experimentally in Ti1-xCrxN by Refs. aivazov_magnetic_1975 ; inumaru_ferromagnetic_2007 at the TiN-rich regime. In Ref. alling_theory_2010 , this system was investigated using first-principles calculations, confirming the experimental findings. The experimental results in Fig. 1 for this system are from Ref. inumaru_ferromagnetic_2007 . When conducting the ground-state search, the chromium moments do not end up completely ferromagnetically aligned in any of the cases $x=0.125,0.25,0.375$, or $0.5$. Judging from the total magnetic moment of the supercell, the average magnetic moment is 1.84, 0.79, 0.55, and 1.12 $\mu_{B}$/chromium atom for the respective $x$-values, reflecting directional variations but no change in magnitude. In the ferromagnetic case, the average chromium moment has a magnitude of approximately $2.2$ $\mu_{B}$ for all $x$-values. The energies of the ground-state search are compared to a ferromagnetic calculation, and in all cases the ground-state search results in a slightly lower energy. The largest difference is for $x=0.5$, where the ferromagnetic calculation results in an energy 13 meV/chromium atom higher than the ground-state search.
@@
In the workflow developed in our previous work, the magnetic moments to be constrained in the DLM run are determined based on the average magnetic moment magnitude of the specific species obtained in the ground-state search. If the average magnetic moment magnitude is larger than 0.75 $\mu_{B}$ in the ferromagnetic ground state, the magnetic moments are constrained; otherwise they are free to change size and direction in the DLM calculations. These atoms are essentially treated on an equal footing with nonmagnetic atoms. Following this strictly in some cases is, however, not possible. For Fe1-xCrx, the chromium moments are on average approximately $1.3$ $\mu_{B}$ when $x=0.125$ and should therefore be constrained in the DLM run. These moments are, despite this, not constrained to avoid convergence problems, which is discussed in more detail in Sec. V. This is also the case for the vanadium moments in Fe1-xVx, where, for all tested $x$-values, these moments are larger than $0.75$ $\mu_{B}$ but are never constrained in the DLM calculations.
@@
The fcc Co1-xAlx system is also investigated using our model. This system is separated from the others as an example of a case with more itinerant type of magnetism to illustrate when our model is not expected to perform as well as in the above examples. This is confirmed in Fig. 2. The procedure of increasing the constraining parameter $\lambda$ is slightly adjusted for this alloy system due to a higher sensitivity in the electronic convergence of the self-consistency cycle. After $\lambda=5$ the $\lambda$ parameter is set to 7. In the case of 25% aluminum, a satisfying level of electronic convergence is not reached for $\lambda=10$. Therefore, Fig. 2 shows the results of $\lambda=7$ for all percentages of aluminum. The penalty energy in these cases is approximately $0.2$ meV/magnetic atom, which we consider an acceptable level of accuracy. The experimental values for Co1-xAlx are from Ref. mcalister_-co_1989 . In Fig. 3a, the magnetic moment sizes of the cobalt atoms are shown for both Co0.9Al0.1 and Co0.9Fe0.1. The moment sizes are given in the DLM state and in the magnetic ground state for comparison. In Fig. 3b, the magnetic moment sizes of iron and aluminum are given for the same systems.
@@
Figure 2: Predicted and experimental Curie temperatures of fcc Co1-xAlx as a function of the fractional substitution of cobalt with aluminum. The experimental Curie temperatures are from Ref. mcalister_-co_1989 .